OpenStudy (anonymous):

GUYS, I'M FANNING AND MEDALIN !!!! PLS HELP!! DESPERATE HERE !!! Solve the following system of equations algebraically. \begin{cases} y = 5x^{2}+30x+41 \\ y = 3x^{2}+24x+46 \end{cases}

OpenStudy (anonymous):

Let y=y 5x^2 +30x +41= 3x^2 +24x +46 subtract 3x^2 , 24x , 46 over to the left side 2x^2 +6x -5 solve this using quadratic formula

OpenStudy (anonymous):

OMG, I LUV YOU SO MUCH <333333 LIFESAVER ;)

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a}\] \[\frac{ -6 \pm \sqrt{36-4(2)(-5)} }{ 2(2)}\] \[\frac{ -6 \pm \sqrt{76} }{ 4}\] \[\frac{ -6 \pm 2\sqrt{19} }{ 4}\] \[\frac{ -3 \pm \sqrt{19} }{ 2}\]

OpenStudy (anonymous):

Anytime :)

OpenStudy (nincompoop):

http://www.ck12.org/algebra/